Project

General

Profile

NDOS CommissionMtng 20130114

Attending: Jarek, Jan, Susan, Nick, Domminick, Gavin, Peter, Steve, Leon, Xuebing

UPDATES:
-DataCheck Update
Jan 3 for 5 days

Stopped for 11:27 - IOC No cooling info for 3 runs - DCM 1-2 channel 34 3-1- 56 shut before and after DSO scan
14418

Jan 10- two RCstart runs for 14419 - run was started on run control machine.
Single empty event in subrun zero is a common occurrence

14421 - 3-1 56 shows up as not instrumented; 1-1 38 shut off
14422 - subrun 3 went for more than 24 hrs Peter gave a tour: paused run, incremented on resume
14423 - Running now

-Installation Plans for the week
MC Decommissioning on hold

-DAQ Update and Plans for the week
Abbey throttled back the trigger debug messages in the Message viewer
Peter wants to shakedown the far detector more
DDT AppMan - Will discuss at the DDT meeting.
- Peter requests a sketch of how DDT plays with DAQ

-DCS Update and Plans for the week
DCSApp manager - Will live on some large partition number (shouldn't interfere)

-AOB

Jarek - UMN connected with VNC last week on desktops 1 and 2. ask for Control specs and requirement
- Asks if they can take control of NDOS
- Andrew: There is no video conferencing solution, Jarek spoke to Andrew about hardware requirements

Mat's last meeting as RC.